A mechanic has 343 gallons of gasoline and 18 gallons of oil to make gas/oil mixtures. He wants one mixture to be 6% oil and the other mixture to be 2.5% oil. If he wants to use all of the gas and oil, how many gallons of gas and oil are in each of the resulting mixtures?
Answer by ankor@dixie-net.com(22740)

A mechanic has 343 gallons of gasoline and 18 gallons of oil to make gas/oil mixtures.
He wants one mixture to be 6% oil and the other mixture to be 2.5% oil.
If he wants to use all of the gas and oil, how many gallons of gas and oil are
in each of the resulting mixtures?
:
The total amt of the two mixtures: 343+18 = 361
:
Let x = amt of the 6% mixture
then
(361-x) = amt of the 2.5% mixtures
:
Write an amt of oil equation
.06x + .025(361-x) = 18
.06x + 9.025 - .025x = 18
.06x - .025x = 18 - 9.025
.035x = 8.975
x = 9.425/.035
x = 256.4 gal for the 6% mixture
and
361 - 256.4 = 104.6 gal for the 2.5% mixture
:
Find the amt of oil in each
.06(256.4) = 15.4 gal of oil in the 6% mixture
.025(104.6) = 2.6 gal of oil in the 2.5% mixture
--------------------
Check total = 18 gal
:
Find the amt of gas in each mixture
.94(256.4) = 241 gal of gas in the 6% mixture
.975(104.6) = 102 gal of gas in the 2.5% mixture
-----------------
Check gas total: 343 gal